OpenStudy (anonymous):

SIMPLE QUESTION: Solve. Check. Explain. :D 2x - 7y = 3 -2x + y = -9 Sytems of equations: Addition/Elimination method (I'm repeatedly getting this question wrong..)

OpenStudy (anonymous):

Add the two equations together, what do you get?

OpenStudy (anonymous):

I'm having trouble adding the 2 equations together.. because x and y both end up being 0

OpenStudy (anonymous):

2x+(-2x)=0 but -7y+y=-6y

OpenStudy (anonymous):

No.. I mean I got this when I added the two equations: 0x + 0y = 6 How did you get -7y+y=-6y

OpenStudy (anonymous):

ok, add the x together, the y together and the constants together the x will give you 0 -7y+y and the coefficients -7+1=-6 so -6y=-6

OpenStudy (anonymous):

OH. Oh my god.. How did I miss that..?

OpenStudy (anonymous):

I dunno =p Should be easy now though

OpenStudy (anonymous):

Can you solve it anyways..? Well..actually. Hang on..let me solve this..and tell me whether or not my answer is right..

OpenStudy (anonymous):

y = -1 and x = -4 That's what I got..

OpenStudy (anonymous):

SORRY. x = 4

OpenStudy (anonymous):

well from -6y=-6 y=-6/-6 y=1 now plug that back in to find x